OpenStudy (anonymous):

What is the axis of symmetry for f(x) = −5x2 − 20x − 10? x = 6 x = −6 x = 2 x = −2

OpenStudy (campbell_st):

the axis of symmetry of a parabola \[y = ax^2 + bx + c\] is found using \[x = \frac{-b}{2a}\] in your question b = -20 and a = -5 just substitute and evaluate